# https://www.vrvis.at

### v1

@@ 3.2 Hierarchical Correspondence Matching
We initialize our model by filling $M$ with $m$ labeled samples consisting of $(x,z)_{j}$ as well as their respective feature pyramids $(h)_{j}^{l}$, where $j=1,...,m$ and $l=1,...,n$. Leaf nodes in level $l=0$ refer to memory labels $z_{j}$. Next, we define connectivity kernels which define children-parent relations across different levels. Kernels remain fixed for every sample and for every level. For more details, see Appendix 0.A.
@@
We formulate the transductive forward pass of our method in Algorithm 1. First, a query $x_{q}$ is encoded via $F$ to obtain $(h)_{q}^{l}$. Then, $G$ takes $(h)_{q}^{l}$ together with memory feature pyramids $(h)_{j}^{l}$ and their respective labels $z_{j}$ as inputs. The overall objective is searching for both globally and locally similar nodes in $M$. Starting from $l=n$, the search objective in each level is first comparing queries $h_{q}^{l}$ with keys $h_{j}^{l}$, followed by retrieving the children of the top-$k$ most similar memory nodes. We use the cosine similarity between nodes’ hidden features as our similarity measure. The accumulated similarity $s_{acc}^{l}$ between two nodes at level $l$ is defined as the product of their similarities $s^{l}$ at level $l$ and the similarities $s_{k,par}$ of their respective parents across all higher levels:
$s_{acc}^{l}=\prod_{u=l}^{n}sim(h_{q}^{u},h_{j}^{u})=\prod_{u=l}^{n}\frac{h_{q}^{u}\cdot h_{j}^{u}}{\lVert h_{q}^{u}\rVert\ \lVert h_{j}^{u}\rVert}$ (1)
For each of the $p^{l}$ query nodes, we use $s_{acc}^{l}$ to identify the top-$k$ most similar nodes in level $l$ and retrieve their respective children in level $l-1$. This implies that the top-$k$ most similar nodes in level $l$ become parent nodes for level $l-1$. Children nodes in level $l-1$ are required for processing level $l-1$, thereby forming a hierarchy across levels. The hyperparameter $\phi\in(0.0,1.0]$ controls the number $k$ of top-$k$ nodes to retain for every query node in each level during search. Effectively, $\phi$ is the rate of reduction for $k$ in each level. A variable number of nodes can be retained in different memory samples.
Considering the data dimension $dim$ and the grid resolutions $res^{l}$, exhaustive search would result in a computational complexity of $\mathcal{O}((\prod_{dim}res_{dim}^{l})^{2})$ for level $l$. In contrast, our approach performs search locally and has a computational complexity of only $\mathcal{O}(n_{ch}\cdot(\prod_{dim}res_{dim}^{l}))$ with $n_{ch}$ being the number of child nodes in level $l+1$. Also, since search is performed independently for every query node, our approach can be highly parallelized in practice and is thus scalable to high resolutions and high-dimensional data. Moreover, since $h^{l}$ are strongly correlated across levels, there is a high chance that similarities to query nodes are also strongly correlated across levels. In other words, searching locally pre-filters nodes that are unlikely to be among the top-$k$ most similar nodes as search progresses to the next level.
@@
Memory search results in a dense correspondence map between a query $x_{q}$ and, for every query node, their top-$k$ most similar memory nodes $i_{k,par}$. We use these matches to retrieve task-specific labels $z_{k}$ from $M$. To obtain a raw prediction $y_{q,raw}$, we use the softmax-normalized similarities $s_{k,par}$ of all matches to obtain a weighted average of retrieved labels.
3.3 Intra-Query Transduction via Message Passing
@@
Up until this point we did not exploit any intra-query correlations such as applying geometric priors or enforcing smoothness of predictions. Strong local correlations are common in grid-based data and should be exploited if possible. The simplest variant of intra-query transduction is Gaussian blurring, where local smoothness over neighboring nodes is enforced. However, blurring is not content-aware and can harm performance at certain regions, e.g., at sharp corners or small objects. We therefore propose a content-aware, parameter-free message passing (MP) approach to account for local correlations within the query and achieve pixel-exact predictions. We re-use our hierarchical search without memory content to find the top-$\kappa$ nearest neighbor nodes within the query sample itself. Similarity scores between a query node $i$ and its nearest neighbors $j$ are referred to as edges $e_{ij}$. We then perform MP to refine raw predictions $y_{q,raw}$. Considering a query node state $y_{q,i}$, we define the equations for the edges $e_{ij}$, the aggregated message $\hat{y}_{q,i}$, and the node update from step $t\rightarrow t+1$ as
@@
The hyperparameter $\lambda\in(0.0,1.0]$ controls the overall strength of MP per step. We apply Eq. 3 and Eq. 4 repeatedly for all query nodes in level $l=0$ until convergence [113], i.e., when $||y_{q,i}^{t+1}-{y}_{q,i}^{t}||\rightarrow 0$. The resulting prediction $y_{q}$ then contains less artifacts and is spatially more consistent compared to $y_{q,raw}$.
3.4 Test-Time Augmentation
Similar to recent work [36], we perform test-time augmentation (TTA) to resolve small errors. We apply multiple forward passes with down-scaled and reflection-padded variants of $x_{q}$. Registration of down-scaled variants with the original variant is achieved by up-scaling and center-cropping the respective predictions. We obtain a final prediction $y_{q}$ by averaging over all passes.
3.5 Fast Learning
PARMESAN is designed to learn fast in the sense that it does not require training of learnable parameters $\theta$ to perform complex tasks such as CL or FSL. In contrast to other approaches, our method only requires memory consolidation, i.e., saving examples in $M$ and extracting their respective feature pyramids $(h)_{j}^{l}$ in a single forward pass using $F$ with pre-trained, frozen $\theta$. For a given hardware setup, learning speed can be split up into required wall-clock times $\tau_{p}$ and $\tau_{l}$ for pre-training and learning, respectively. We define $\tau_{p}$ on a per-dataset basis and $\tau_{l}$ on a per-sample basis. The total required wall-clock time to learn $n$ training samples is then $T=\tau_{p}+n\cdot\tau_{l}$.
@@ 4 Experiments and Results
We demonstrate the capabilities of our method in three groups of experiments:
A0 – A18:
ablation studies for semantic segmentation and monocular depth estimation to demonstrate flexibility (Sec. 4.1, Tab. 1, and Tab. 2),
B0 – B7:
CL experiments for studying knowledge retention properties of our method (Sec. 4.2, Tab. 3, and Tab. 4),
C0 – C7:
FSL experiments to investigate data-efficiency (Sec. 4.3 and Tab. 5).
More details regarding applied models, training and evaluation, as well as qualitative results for all experiments are provided in Sec. 0.B.2.
Our main goal is to improve flexibility and adaptation capabilities of DL models while simplifying and speeding up learning for CL and FSL. Problem-specific methods are expected to outperform more general methods such as ours [57, 104, 101].
Datasets: We use the Cityscapes (CITY) benchmark dataset [7] for semantic segmentation and monocular depth estimation. Semantic segmentation masks comprise fine and coarse labels, whereas disparity maps are provided for depth estimation. We also use the JSRT Chest X-ray dataset [84] for semantic segmentation. We refer to Sec. 0.B.1 for further details.
Method: We employ a U-Net [77] variant as our default DL model and feature extractor. For experiments A11 and A14 we use ConvNeXt Tiny [50] with $\theta$ trained on ADE20K [112]. For all experiments that require $M$, we select memory samples randomly from the training split while keeping the random generation process fixed to ensure comparability. For PARMESAN, we set $m~{}=~{}256$, $\phi=0.5$, apply message passing (MP) and test-time augmentation (TTA), and use both pre-trained U-Net encoder and decoder for feature extraction unless stated otherwise. We use Gaussian blurring with $\sigma=5$ pixels in experiment A2. CL and FSL are performed by memory consolidation of training samples. To obtain memory samples for the privacy-preserving experiment A6, we additionally apply data distillation via linear mixup of two random training samples [109, 43, 83, 70].
Baselines: Joint training (JOINT) refers to training on all available training data, thus representing the default supervised learning setup and the upper bound for CL [97]. Fine-tuning (FT) is a simple, yet strong baseline for FSL [10]. Due to catastrophic forgetting, however, FT is regarded as the lower bound for CL [97]. We further employ memory-based approaches such as classical replay (REPLAY) [82, 71, 51, 76] and greedy-sampling dumb learning (GDUMB) [66] as additional baselines for CL. With GDUMB, training is performed from scratch for each individual CL step, thereby exclusively using memory samples. For consistent and fair comparisons to our method, we implement all baselines and adjust method-specific hyperparameters to our data setup.
Evaluation: We use the mean Intersection over Union of classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$) as evaluation metrics for semantic segmentation and Root-Mean-Square Error (RMSE) in meters for depth estimation. We report best performances on the CITY validation set and on the JSRT test set, since labels for the CITY test set are not publicly available.
@@ 4.1 Ablation Studies
We perform ablation studies of various components to analyze the impact on learning speed and predictive performance, as well as demonstrating the flexibility of our method with different architectures of $F$. We summarize our results in Tab. 1 (segmentation) and Tab. 2 (depth) and visualize predictions in Fig. 2 (CITY, segmentation), Fig. 3 (CITY, depth), and Fig. 4 (JSRT, segmentation).
Learning with our method can be done in less than 0.1 seconds per sample (see column $\tau_{l}$ in Tab. 1 and Tab. 2), thereby mitigating the demand for additional computational costs when attaching PARMESAN to existing models. Our method outperforms JOINT on CITY when trained on coarse labels (A9, A10) and evaluated on fine labels. As expected, our method under-performs JOINT when using fine labels (A0, A1), on JSRT (A12, A13, A14), and for depth estimation (A15, A16). Using only the encoder (A7) or using an entirely untrained, randomly initialized model (A8) yields promising results already, indicating that our transduction module $G$ imposes useful inductive biases. PARMESAN can be successfully used in combination with a ConvNeXt feature extractor on both CITY and JSRT when trained on ADE20K [50] (A11, A14). MP and TTA (A1, A3, A4, A16, A17) considerably improve performance over raw predictions (A5, A18). As expected, content-aware MP (A1) also outperforms non-content-aware Gaussian blurring (A2). Comparing panels A1 and A5 in Fig. 2, MP and TTA succeed in reducing small artifacts while retaining sharp corners of predicted segments. In contrast, blurring typically comes at the cost of missing small segments and sharp corners. Privacy-preservation by data distillation (A6) slightly decreases performance compared to the non-private setup (A1).
Figure 2: Model predictions and analysis of PARMESAN. Left: We study top-$1$ nearest neighbors from experiment A1. The idx panel visualizes indices of retrieved nearest neighbor nodes from $M$. Similar colors indicate that nodes are retrieved from close-by memory locations, most likely from within the same samples. The sim panel visualizes the respective similarity scores, where bright regions refer to high similarities. We also show 3 query pixels and their matches in $M$, indicating both global and local semantic correspondence. Right: PARMESAN can be applied flexibly (A5, A11) and perform complex tasks like CL (B5, B7) and FSL (C3, C7) via memory consolidation.
Figure 3: Model predictions for monocular depth estimation experiments.
Figure 4: Model predictions for semantic segmentation experiments using JSRT data.
@@
Table 1: Results for ablation studies showing semantic segmentation performance for classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$). Column “$F$ pre-train/train“ indicates if the method uses a pre-trained feature extractor $F$ and if any additional training of learnable parameters $\theta$ is required. $\tau_{p}$ and $\tau_{l}$ are wall-clock times for pre-training and learning, respectively. Experiment IDs marked with * use ConvNeXt pre-trained on ADE20K [50] for feature extraction.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
A0 CITY, fine (2975) JOINT no / yes 0 3.72 82.5 87.4
A1 CITY, fine (256) ours A0 / no 3.08 0.09 80.0 85.5
A2 CITY, fine (256) ours, no MP, blur A0 / no 3.08 0.09 79.3 84.8
A3 CITY, fine (256) ours, no TTA A0 / no 3.08 0.09 77.7 83.5
A4 CITY, fine (256) ours, no MP A0 / no 3.08 0.09 77.9 83.2
A5 CITY, fine (256) ours, no MP, no TTA A0 / no 3.08 0.09 75.7 81.6
A6 CITY, fine (256) ours, private A0 / no 3.08 0.10 77.6 83.2
A7 CITY, fine (256) ours A0 (E only) / no 3.08 0.08 72.8 79.1
A8 CITY, fine (256) ours no / no 0 0.09 56.1 63.5
A9 CITY, coarse (19998) JOINT no / yes 0 0.43 49.6 50.6
A10 CITY, coarse (256) ours A9 / no 2.41 0.09 50.7 52.5
A11* CITY, fine (256) ours E only / no - 0.09 74.5 81.4
A12 JSRT (199) JOINT no / yes 0 23.39 93.3 -
A13 JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ A12 / no 1.29 0.02 91.9 -
A14* JSRT (199) ours, $m\!=\!199$, $\phi\!=\!0.8$ E only / no - 0.02 87.2 -
Table 2: Results for ablation studies showing depth estimation performance.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] RMSE [m]
A15 CITY, disparity (2975) JOINT no / yes 0 6.71 10.5
A16 CITY, disparity (256) ours A15 / no 5.55 0.03 11.9
A17 CITY, disparity (256) ours, no MP A15 / no 5.55 0.03 12.5
A18 CITY, disparity (256) ours, no MP, no TTA A15 / no 5.55 0.03 13.5

### v2

@@ 3.2 Correspondence Matching and Intra-Query Transduction
We initialize our model by filling $M$ with $m$ labeled samples $(x,z,(h)^{l})_{j}$, where $j=1,...,m$ and $l=1,...,n$. Leaf nodes in $l=0$ refer to $z_{j}$. Next, we define connectivity kernels which define children-parent relations across different levels. Kernels remain fixed for every sample and every level (also see Appendix 0.A).
@@
We formalize the forward pass of our method in Algorithm 1. First, a query $x_{q}$ is encoded by $F$ to get $(h)_{q}^{l}$. Then, $G$ takes $(h)_{q}^{l}$ together with stored $(h)_{j}^{l}$ and $z_{j}$ as inputs. The overall objective is to find globally and locally similar patterns in $M$ w.r.t. $x_{q}$. Traversing levels backwards starting from $l=n$, we compare queries $h_{q}^{l}$ with keys $h_{j}^{l}$, followed by retrieving the children (within $l-1$) of the top-$k$ most similar nodes of $h_{j}^{l}$. We use cosine similarity by default, where the accumulated similarity $s_{acc}^{l}$ between two nodes at level $l$ is the product of their similarities $s^{l}$ at level $l$ and the similarities $s_{k,par}$ of their respective parents across all higher levels:
$s_{acc}^{l}=\prod_{u=l}^{n}sim(h_{q}^{u},h_{j}^{u})=\prod_{u=l}^{n}\frac{h_{q}^{u}\cdot h_{j}^{u}}{\lVert h_{q}^{u}\rVert\ \lVert h_{j}^{u}\rVert}\ .$ (1)
@@
For each of the $p^{l}$ query nodes, we use $s_{acc}^{l}$ for comparisons. The top-$k$ most similar nodes in $l$ become parents for $l-1$. Children in $l-1$ are required for processing $l-1$, thereby forming a hierarchy across levels. The reduction rate $\phi\in(0.0,1.0]$ controls $k$, i.e., how many nodes are retained in each level. A variable number of nodes can be retained in different memory samples. Memory search results in a dense correspondence map between a query $x_{q}$ and, for every query node, their top-$k$ most similar memory nodes $i_{k,par}$. We use these matches to retrieve task-specific labels $z_{k}$ from $M$. We obtain a raw prediction by weighting retrieved labels with softmax-normalized similarities $s_{k,par}$ of all matches.
Considering the data dimension $dim$ and the grid resolutions $res^{l}$, exhaustive search would result in a computational complexity of $\mathcal{O}((\prod_{dim}res_{dim}^{l})^{2})$ for level $l$. In contrast, our method performs local search and has a complexity of only $\mathcal{O}(n_{ch}\cdot(\prod_{dim}res_{dim}^{l}))$ with $n_{ch}$ being the number of children of nodes in $l+1$. Search is performed independently and in parallel for every query node, making our approach scalable to high resolutions. Moreover, since $h^{l}$ are strongly correlated across levels, there is a high chance that similarities to query nodes are also strongly correlated across levels. In other words, searching locally pre-filters nodes that are unlikely to be among the top-$k$ most similar nodes in the next level anyway.
@@
Up until this point, no intra-query correlations had been exploited by using geometric or smoothness priors. Strong local correlations are common in grid-based data and should be exploited if possible. We therefore propose a content-aware, parameter-free message passing (MP) approach to account for local correlations within the query and achieve pixel-exact predictions. Re-using our hierarchical search without memory content, top-$\kappa$ nearest neighbor nodes within the query itself are found. Similarity scores between a query node $i$ and its nearest neighbors $j$ are referred to as edges $e_{ij}$. MP is then perfomed to refine raw predictions $y_{q,raw}$. Considering a query node state $y_{q,i}$, we define the equations for the edges $e_{ij}$, the aggregated message $\hat{y}_{q,i}$, and the node update from step $t\rightarrow t+1$ as
@@
The hyperparameter $\lambda\in(0.0,1.0]$ controls the overall strength of MP per step. We apply Equations 3 and 4 repeatedly for all query nodes in level $l=0$ until convergence [100], i.e., when $||y_{q,i}^{t+1}-{y}_{q,i}^{t}||\rightarrow 0$. The resulting prediction $y_{q}$ then contains fewer artifacts and is spatially more consistent compared to $y_{q,raw}$.
3.3 Memory-Efficiency via Sparsity and Novelty
In its basic setup, PARMESAN requires considerable hardware memory resources to operate. We thus introduce sparsity, i.e., a reduction of the number of patterns per stored sample, to become exponentially more memory efficient (see Fig. 2). Sparsity effectively reduces spatial resolutions of stored samples and is applied to both $(h)_{j}^{l}$ and $z_{j}$ after feature extraction of full-resolution inputs $x_{j}$. Starting from $l=1$, the number of sparse levels $n_{sp}$ defines the overall degree of sparsity. Naive sparsity can be achieved by random subsampling of nodes in levels $l=1,...,n_{sp}$. However, we apply local novelty search to retain high diversity in memory contents. Starting from $l=1$, we keep the most novel node in every $2^{dim}$ patch such that sparsified resolution $res_{sp,dim}^{1}=res_{dim}^{2}$. After applying sparsity to $l=2$ such that $res_{sp,dim}^{2}=res_{dim}^{3}$, we discard previously kept nodes in $l=1$ such that $res_{sp,dim}^{1}=res_{sp,dim}^{2}=res_{dim}^{3}$. We repeat this procedure for every sparse level and independently for all memory samples.
Figure 2: Memory sample feature pyramid with $n_{sp}=2$ sparse levels obtained via iterative, local novelty search (indicated by circles). Filled nodes are kept in $M$.
@@ 4 Experiments and Results
We demonstrate the capabilities of our method in ablation studies (Section 4.1) and CL experiments (Section 4.2). For ablations, we aim to demonstrate flexibility and efficiency while simplifying and speeding up learning. For CL, we study knowledge retention and overall performance. More details regarding applied models, training and evaluation, and additional results are provided in Section 0.B.1 and 0.B.2.
Datasets: We use the Cityscapes (CITY) dataset [8] for semantic segmentation and monocular depth estimation. Segmentation masks comprise fine and coarse labels. Disparity maps are given for depth estimation. We also use the JSRT Chest X-ray dataset [76] for semantic segmentation.
Method: We employ a U-Net variant as our default DL model and feature extractor $F$. Some experiments use a ConvNeXt Tiny [45] encoder pre-trained on ADE20K [99]. If $M$ is used, we select memory samples randomly from the training split while keeping the random generation process fixed across all experiments. For PARMESAN, we set $m\!=\!2975$, $\phi\!=\!0.5$, $n_{sp}\!=\!3$ ($m\!=\!199$, $\phi\!=\!0.8$, $n_{sp}\!=\!3$ for JSRT), apply message passing (MP), and take decoder features from $F$ unless stated otherwise. CL is performed by memory consolidation of training samples. We also use test-time augmentation (TTA) to boost performance [31].
Baselines: We select baselines with the premise to compare to representative and well-established methods in the field. Joint training (JOINT) refers to training on all training data, thus representing the default supervised learning setup and the upper bound for CL [88]. Due to catastrophic forgetting, fine-tuning (FT) is regarded as the lower bound for CL [88]. We employ classical replay (REPLAY) [75, 63, 46, 68], greedy-sampling dumb learning (GDUMB) [60], and Modeling the Background (MiB) [5] as additional CL baselines. Training via GDUMB is performed from scratch for each individual CL step, thereby exclusively using memory samples. For fair and consistent comparisons, we implement all baselines and adjust method-specific hyperparameters to our data setup.
Evaluation: We use the mean Intersection over Union of classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$) as evaluation metrics for semantic segmentation and Root-Mean-Square Error (RMSE) in meters for depth estimation. We report best performances on the CITY validation set and on the JSRT test set, since labels for the CITY test set are not publicly available. For a given hardware setup, the learning speed $\tau_{l}$ is the average wall-clock time required to learn a training sample. The total learning time for $n$ samples is then $T=n\cdot\tau_{l}$.
@@ 4.1 Ablation Studies
We perform ablation studies of various components to analyze the impact on learning speed and predictive performance, as well as demonstrating the flexibility of our method with different architectures of $F$ (see Table 1).
Performance of our method is on-par with JOINT (A0 / A1, A9 / A10, A12 / A13) while greatly enhancing flexibility and learning speed. Learning with PARMESAN can be done in $\tau_{l}\simeq 0.01$ seconds per sample, thereby mitigating the demand for additional computational costs when attached to existing models. Using only the encoder (A4) or even random weights (A5) yields promising results already, indicating that $G$ imposes useful inductive biases. PARMESAN can be successfully used for transfer learning when using a publicly available ConvNeXt pre-trained on ADE20K (A8, A11) or pre-training on coarse labels (A7). MP and TTA (A1 / A2 / A3) considerably improve raw predictions and succeed in reducing small artifacts while retaining sharp corners of predicted segments (see Fig. 3).
Figure 3: Model predictions and analysis of PARMESAN. Left: We study top-$1$ nearest neighbors from experiment A1. The idx panel visualizes indices of retrieved nearest neighbor nodes from $M$. Similar colors indicate labels retrieved from close-by memory locations, most likely from within same samples. The sim panel visualizes $s_{acc}^{1}$, where bright regions refer to high similarities. We also show 3 query pixels and their matches in $M$, indicating both global and local semantic correspondence. Right: Flexible use of PARMESAN in various setups. Bottom: Depth estimation on CITY and semantic segmentation on JSRT.
@@
In Fig. 4, we demonstrate that our novelty-sparsity approach leads to exponentially lower memory requirements while achieving robust predictive performance.
Table 1: Results for ablation studies, demonstrating the capabilities of PARMESAN. For semantic segmentation, we report performance for classes ($\text{mIoU}_{\text{cl}}$) and categories ($\text{mIoU}_{\text{cat}}$). “$\theta$ pre-train” and “$\theta$ train” indicate the usage of a pre-trained feature extractor $F$ and if any additional training of learnable parameters $\theta$ is required. $\tau_{l}$ is the average wall-clock time for learning a sample.
ID Data (#train-samples) Method & Setup $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
A0 CITY, fine (2975) JOINT 3.72 82.5 87.4
A1 CITY, fine (2975) ours A0 0.012 82.1 87.0
A2 CITY, fine (2975) ours, no MP A0 0.012 81.4 86.4
A3 CITY, fine (2975) ours, no TTA A0 0.012 81.2 86.3
A4 CITY, fine (2975) ours, E only A0 0.009 75.6 81.6
A5 CITY, fine (2975) ours, random $\theta$ 0.012 59.6 66.3
A6 CITY, coarse (19998) JOINT 0.43 49.0 50.3
A7 CITY, fine (2975) ours A6 0.012 76.9 82.4
A8 CITY, fine (2975) ours, ConvNeXt (E) [45] 0.021 78.2 84.3
A9 JSRT (199) JOINT 23.4 93.9 -
A10 JSRT (199) ours A9 0.009 92.6 -
A11 JSRT (199) ours, ConvNeXt (E) [45] 0.011 84.7 -
ID Data (#train-samples) Method & Setup $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] RMSE [m]
A12 CITY, disparity (2975) JOINT 6.71 10.5
A13 CITY, disparity (2975) ours A12 0.01 11.5
Figure 4: Performance w.r.t. memory size on CITY. We control memory size by introducing sparse levels, leading to exponentially lower memory requirements and stable performance. Relevant contents are retained using local novelty search.